# The flexible Gumbel distribution: A new model for inference about the mode

### v2

@@ 7 Discussion
The mode had been an overlooked location parameter in statistical inference until recently when the statistics community witnessed a revived interest in modal regression among statisticians [Chen, 2018, Chacón, 2020, Feng et al., 2020, Xu et al., 2020, Ullah et al., 2021, Wang and Li, 2021, Xiang and Yao, 2022]. Historically, statistical inference for the mode have been mostly developed under the nonparametric framework for reasons we point out in Section 1. Existing semiparametric methods for modal regression only introduce parametric ingredients in the regression function, i.e., the conditional mode of the response, with the mode-zero error distribution left in a nonparametric form [Yao and Li, 2013, Liu et al., 2013, Zhang et al., 2013, Yang and Yang, 2014, Zhao et al., 2014, Krief, 2017, Tian et al., 2017, Li and Huang, 2019]. The few recently proposed parametric modal regression models all impose stringent parametric assumptions on the error distribution [Bourguignon et al., 2020, Zhou and Huang, 2020, 2022]. Our proposed flexible Gumbel distribution greatly alleviates concerns contributing to data scientists’ reluctance to adopt a parametric framework when drawing inference for the mode. This new distribution is a heterogeneous mixture in the sense that the two components in the mixture belong to different Gumbel distribution families, which is a feature that shields it from the non-identifiability issue most traditional mixture distributions face, such as the normal mixtures. The proposed distribution is indexed by the mode along with shape and scale parameters, and thus is convenient to use to draw inference for the mode while remaining flexible. It is also especially suitable for modeling heavy-tailed data, whether the heaviness in tails is due to extremely large or extremely small observations, or both. These are virtues of FG that cannot be achieved by the popular normal mixture and many other existing mixture distributions.

### v3

@@ 7 Discussion
The mode had been an overlooked location parameter in statistical inference until recently when the statistics community witnessed a revived interest in modal regression among statisticians [Chen, 2018, Chacón, 2020, Feng et al., 2020, Xu et al., 2020, Ullah et al., 2021, Wang and Li, 2021, Xiang and Yao, 2022]. Historically, statistical inference for the mode has been mostly developed under the nonparametric framework for reasons we point out in Section 1. Existing semiparametric methods for modal regression only introduce parametric ingredients in the regression function, i.e., the conditional mode of the response, with the mode-zero error distribution left in a nonparametric form [Yao and Li, 2013, Liu et al., 2013, Zhang et al., 2013, Yang and Yang, 2014, Zhao et al., 2014, Krief, 2017, Tian et al., 2017, Li and Huang, 2019]. The few recently proposed parametric modal regression models all impose stringent parametric assumptions on the error distribution [Bourguignon et al., 2020, Zhou and Huang, 2020, 2022]. Our proposed flexible Gumbel distribution greatly alleviates concerns contributing to data scientists’ reluctance to adopt a parametric framework when drawing inferences for the mode. This new distribution is a heterogeneous mixture in the sense that the two components in the mixture belong to different Gumbel distribution families, which is a feature that shields it from the non-identifiability issue most traditional mixture distributions face, such as the normal mixtures. The proposed distribution is indexed by the mode along with shape and scale parameters, and thus is convenient to use to draw inferences for the mode while remaining flexible. It is also especially suitable for modeling heavy-tailed data, whether the heaviness in tails is due to extremely large or extremely small observations, or both. These are virtues of FG that cannot be achieved by the popular normal mixture and many other existing mixture distributions.
@@
Appendix A Derivation of the third central moment of FG in (4)
For any finite mixture distribution, its higher-order central moments can be expressed using the binomial formula (See Equation (1.22) in [Frühwirth-Schnatter, 2006]). The FG distribution is a mixture distribution with two components: a right-skewed Gumbel distribution and a left-skewed Gumbel distribution. Let $Y\sim\operatorname{FG}\left(\theta=0,\sigma_{1},\sigma_{2},w\right)$, $Y_{1}\sim\text{right-skewed Gumbel}\left(\theta=0,\sigma_{1}\right)$, and $Y_{2}\sim\text{left-skewed Gumbel}\left(\theta=0,\sigma_{2}\right)$. Its $j$-th finite moment can be expressed as
$\mathbb{E}\left\{\left(Y-\mu_{y}\right)^{j}\right\}$ $=w\mathbb{E}\left\{\left(Y_{1}-\mu_{1}+\mu_{1}-\mu_{y}\right)^{j}\right\}+\bar{w}\mathbb{E}\left\{\left(Y_{2}-\mu_{2}+\mu_{2}-\mu_{y}\right)^{j}\right\}$ (9)
$=\sum_{k=0}^{j}w\mathbb{E}\left\{\left(\begin{array}[]{l}j\\ k\end{array}\right)\left(Y_{1}-\mu_{1}\right)^{k}\left(\mu_{1}-\mu_{y}\right)^{j-k}\right\}$
$\quad+\bar{w}\mathbb{E}\left\{\left(\begin{array}[]{l}j\\ k\end{array}\right)\left(Y_{2}-\mu_{2}\right)^{k}\left(\mu_{2}-\mu_{y}\right)^{j-k}\right\},$
where $\mu_{y}$, $\mu_{1}$, and $\mu_{2}$ are the expectations of $Y$, $Y_{1}$, and $Y_{2}$, respectively. Applying (9) for $j=3$, one obtains the expression in (4).
Appendix B Proof of identifiability of FG in (3)
In our context of two-component mixture distributions of which the cumulative distribution functions are of the form $wF_{1}(x)+\bar{w}F_{2}(x)$, Theorem 1 in Teicher [1961] states that a mixture distribution is identifiable if and only if there exists $y_{1}$ in the support of $F_{1}(y)$ and $y_{2}$ in the support of $F_{2}(y)$ such that
$\left|\begin{array}[]{cc}F_{1}\left(y_{1}\right)&F_{2}\left(y_{1}\right)\\ F_{1}\left(y_{2}\right)&F_{2}\left(y_{2}\right)\end{array}\right|\neq 0,$
that is, the above determinant does not vanish for some $(y_{1},y_{2})$. In what follows, we prove that the FG distribution is identifiable by showing the existence of $(y_{1},y_{2})$ that makes the above determinant non-zero.
Proof.
Recall that the cumulative distribution functions of right-skewed and left-skewed Gumbel distributions are given by
$F_{1}(y)=\exp\left\{-\exp\left(-\frac{y-\theta}{\sigma_{1}}\right)\right\},$
and
$F_{2}\left(y\right)=1-\exp\left\{-\exp\left(\frac{y-\theta}{\sigma_{2}}\right)\right\},$
respectively.
By setting $y_{1}=\theta$, we have
$\left|\begin{array}[]{cc}F_{1}\left(y_{1}\right)&F_{2}\left(y_{1}\right)\\ F_{1}\left(y_{2}\right)&F_{2}\left(y_{2}\right)\end{array}\right|=\left|\begin{array}[]{cc}e^{-1}&1-e^{-1}\\ F_{1}\left(y_{2}\right)&F_{2}\left(y_{2}\right)\end{array}\right|=e^{-1}F_{2}(y_{2})-(1-e^{-1})F_{1}(y_{2}).$ (10)
We next show by contradiction that there exists $y_{2}\in\mathbb{R}$ such that (10) is not equal to zero.
Suppose (10) is equal to zero for all $y_{2}\in\mathbb{R}$, that is,
$F_{2}(y_{2})=(e-1)F_{1}(y_{2}),\text{ for all $y_{2}\in\mathbb{R}$}.$
Taking the limit of both sides of the above equation as $y_{2}\to+\infty$ gives
$\lim_{y_{2}\rightarrow+\infty}F_{2}\left(y_{2}\right)=(e-1)\times\lim_{y_{2}\rightarrow+\infty}F_{1}\left(y_{2}\right),$
which is clearly false since $\lim_{y_{2}\to+\infty}F_{2}(y_{2})=\lim_{y_{2}\to+\infty}F_{1}(y_{2})=1$. Hence, there exists $y_{2}\in\mathbb{R}$ such that (10) is not equal to zero. Denote by $y_{2}^{*}$ such a value, or one of such values if such $y_{2}$ is not unique.
Now that we have found $(y_{1},y_{2})=(\theta,y_{2}^{*})$ such that the aforementioned determinant does not vanish, the FG distribution is identifiable by Theorem 1 in Teicher [1961]. ∎
Appendix C Convergence diagnosis of MCMC
## hydrology example - Section 5mean se_mean   sd     2.5%      25%      50%      75%    97.5%    n_eff Rhatw1         0.61       0 0.14     0.31     0.51     0.63     0.72     0.83 149860.5    1theta     -0.42       0 0.70    -1.62    -0.93    -0.48     0.04     1.04 162508.9    1sigma1     5.54       0 0.65     4.67     5.11     5.40     5.81     7.25 159083.7    1sigma2     5.95       0 1.04     4.58     5.20     5.73     6.49     8.49 175171.4    1## criminology example - Section 6variable    mean  median      sd     mad      q5      q95  rhat ess_bulk ess_tail<chr>      <dbl>   <dbl>   <dbl>   <dbl>   <dbl>    <dbl> <dbl>    <dbl>    <dbl>1 alpha     0.496   0.530   2.72    2.68   -4.00     4.89    1.00  102742.   41503.2 beta[1]  -0.160  -0.162   0.0785  0.0769 -0.286   -0.0292  1.00   35063.   13152.3 beta[2]   0.235   0.232   0.124   0.120   0.0374   0.441   1.00    2074.    1702.4 beta[3]   0.0669  0.0669  0.0144  0.0140  0.0434   0.0904  1.00    9603.    9166.5 scale1    1.95    1.69    2.36    0.214   1.40     2.64    1.01     588.     189.6 scale2   53.6    38.6    74.2    19.3     1.80   133.      1.01     583.     194.